# Towards Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

### v2

@@ B.1 Image classification
For the DenseNet trained on CIFAR-100, we extract a subset of CIFAR’s training set containing 5000 images to create a proper validation set. We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov (Nesterov, 1983; Sutskever et al., 2013), a momentum of 0.9, and $1\times 10^{-4}$ weight decay. We divide the learning rate by ten after 150 and 225 epochs. We use standard augmentations. We apply random crop with a four-pixel padding as well as random horizontal flip. We do not perform model selection and keep the last checkpoint.
@@
For our ViT-S-16 trained on ImageNet, we take batches of 2048 images and train on 8 V100 for 300 epochs with AdamW (Loshchilov & Hutter, 2019) with the $\beta$s equal to 0.9 and 0.999. We start with a linear warmup for 15 epochs, then use a cosine annealing scheduler with $3\times 10^{-3}$ as the starting learning rate. The models are trained with non-adaptive sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022) with $\rho=0.2$. We use a dropout (Srivastava et al., 2014) rate of 0.1 but no attention dropout. We transform the training images with a standard random resized crop to 224$\times$224 pixels using bicubic interpolation and a random horizontal flip. For evaluation, we center-crop the images to this resolution. For ImageNet, we do not perform model selection and keep the last checkpoint. However, we randomly extract a validation set of 50000 images from the training set to perform logit normalisation.
@@
Our ResNet-50 is trained on 8 V100 with stochastic gradient descent for 120 epochs using a batch size of 1024 images. After five epochs of linear warmup, we use a cosine annealing scheduler starting with a learning rate of 0.4 with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We use the same transformation of the images as for the ViT and select the last checkpoint for inference. We use the same validation set as for the ViT for logit normalisation.
B.2 Semantic segmentation
@@
We train a Deeplabv3+ on CityScapes (Cordts et al., 2016) with a ResNet-101 (He et al., 2016) backbone pre-trained on ImageNet (Russakovsky et al., 2015). We use stochastic gradient descent with a base learning rate of 0.01, divided by 10 for the backbone weights, and reduced following the "poly" policy (Liu et al., 2015) with a power of 0.9. The weights are optimised with a momentum of 0.9 and a weight decay of $10^{-4}$. We take a batch size of 12 images and train for 40000 steps. During training, we randomly crop the input images and targets to squares with 768-pixel-long sides. We apply random horizontal flip and colour-jitter with the classical parameters: brightness, contrast, and saturation levels of 0.5. For testing, we use the images at their original resolution and do not perform any test time augmentations. For the RC curves, we randomly sample 5000 predictions per image, extracted prior to the final interpolation to reduce correlations between the predictions. We keep the pixel-wise locations of the samples when changing the level of label smoothing $\alpha$ to ensure fair comparisons.
B.3 Logit Normalisation
@@
Appendix C Analysis of empirical loss (one-hot labels)
@@
which in turn gives the regularisation gradient,
$\frac{\partial\mathcal{L}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}-\mathcal{L}_{\text{CE}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}}{\partial v_{k}}=\alpha\left[\delta_{y\omega_{k}}-1/K\right]~{},$ (20)
which is once again independent of the model output $\boldsymbol{\pi}$. In this case, as the label $y$ has already been sampled, the model is either right  or wrong , giving two different regularisation gradients for the max logit $v_{\text{max}}$,
@@
We can see that the max logit is more strongly regularised/suppressed during training when the prediction is correct, which aligns with the analysis of $\mathcal{L}^{\text{true}}$ and $P_{\text{error}}$ in the main paper. When the model is correct during training, the max logit is suppressed, but when it is incorrect the max logit is not suppressed. Thus the uncertainty ranking of correct ✓ vs incorrect ✗ data samples is degraded, harming SC. This leads us to the same conclusions as in Sec. 4 and also aligns with the behaviour in Fig. 5 (right) where $v_{\text{max}}$ is higher for errors given the value of MSP.
Appendix D Result and proof
@@
As for Eq. 23, please note that using $k=i$, we get the same equation as Eq. 25, although with a strict inequality. We can now sum on the elements of $\boldsymbol{v}$ to get
@@
By taking the inverse, we get
@@
Appendix E Discussions
E.1 $U$ other than MSP
@@
Finally, we also choose to omit various training/architecture-based approaches such as (Ziyin et al., 2019; Geifman & El-Yaniv, 2019; Moon et al., 2020; Zhu et al., 2024), as we aim to focus our investigation solely on understanding the training effects of LS. We note that LS (potentially combined with post-hoc logit normalisation) is considerably simpler to implement than the aforementioned training-based approaches, and has been shown to be stable in many more use-cases and so may be more likely to be chosen by a practitioner. Besides, recent work (Feng et al., 2023) has suggested that MSP applied to the methods in (Ziyin et al., 2019; Geifman & El-Yaniv, 2019) actually performs better than the $U$s proposed in them (reject logit/selection head).
@@
E.2 Aleatoric and epistemic uncertainty
@@
E.3 On the assumption that logits are positive
@@
Result 1 assumes that all elements in the logit vector are $>0$, which is not necessarily true. However, we also find empirically that both $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ tend to be dominated by the largest positive logits. This is intuitive as exponentiating or raising to power $p>1$ will amplify the larger logits. This is shown in Fig. 19, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set ($p=5$ is optimal on the validation data for logit normalisation in this case). Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive. As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to increase uncertainty for higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$.
@@
E.4 Existing benchmarks and training recipes with LS
E.5 Logit normalisation in (Cattelan & Silva, 2024)
Explanations of effectiveness.
The work that introduces logit normalisation (Cattelan & Silva, 2024) is primarily an experimental study, where the focus is on extracting empirical takeaways. However, the authors do discuss potential reasons for the effectiveness of the approach in their Appendix B. They observe that models that are generally (on both ✓ and ✗) more uncertain benefit from logit normalisation and so suggest that logit normalisation alleviates “underconfidence” by reducing the influence of lower-ranked logits on the uncertainty of a prediction (using analysis similar to ours in Appendix D). We believe their explanation is ultimately incomplete as:
@@ E.5 Logit normalisation in (Cattelan & Silva, 2024)
    They do not clearly delineate the definition of over/underconfidence used in model calibration with that used in selective classification. This is an issue since calibration is concerned with absolute marginal (averaged over data samples) properties, whilst selective classification is concerned with relative conditional (per sample) properties. It is possible to be extremely over/underconfident in the calibration sense, whilst being optimal for SC, or very well calibrated and worse at distinguishing correct vs incorrect samples Zhu et al. (2024).
@@
Logit centralisation.
Cattelan & Silva (2024) suggest an additional step of logit centralisation,
$\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}-\mu(\boldsymbol{v})}{\left\lVert\boldsymbol{v}-\mu(\boldsymbol{v})\right\rVert_{p}},\quad\mu(\boldsymbol{v})=\frac{1}{K}\sum_{k}v_{k},$ (28)
which in practice brings most of the lower valued logits close to zero if they are not already. They note that for the majority of models this step is unnecessary, and we also find it unnecessary in our experiments so we omit it for the sake of simplicity. We note that according to the analysis in Sec. 5.1 and Sec. E.3, for logit normalisation to be effective, $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ should be dominated by positive logits, which centralisation seems to empirically ensure by making most of the smallest logits close to zero (Cattelan & Silva, 2024).
Appendix F Additional segmentation results
In this section, we complete the picture by providing additional semantic segmentation results to Figures 1 and 8. As in the main paper, the following predictions are performed by a ResNet-101-based DeepLab-v3+ trained on Cityscapes. In these figures, we provide the rank of all the predictions – not only errors – to show the difference between the model’s behavior on its errors but also correct predictions. The rank of the correct predictions ✓ is slightly greyed out compared to the rank of the errors ✗. Figure 21 presents the full ranks corresponding to the scene used in the main paper. We will provide other examples to the readers upon request. Later, we will make our models available to ease the reproducibility of our work.
@@ Appendix F Additional segmentation results
Figure 22 shows that the cross-entropy-based model predicts label maps in which pixel uncertainty ranks are smoother than those of the label-smoothing-based one. The label-smoothing model exhibits large areas with low uncertainty rankings, which suddenly decreases nearby the boundary of objects. However, the boundaries are sometimes wrongly predicted, leading to (rank-wise) highly confident errors ✗. The pixels corresponding to the boundaries of objects, where naturally the probability of error is higher and which are most sensitive to label errors, have higher confidence (lower uncertainty ranks) according to the LS-based model than the CE-based model. This explains the high-confidence errors ✗ in the misclassified boundaries of several objects, either in Figs. 1 and 22, and aligns with the imbalanced regularisation of Eq. 15.
We also provide Figure 23 to show that the logit normalisation method is not always successful in solving the ranking of the predictions. In this figure, we see that a large part of the misclassified pixels (lower left) keeps low uncertainty ranks either using the CE-based model or the LS-based model while using logit normalisation. However, there are some improvements on the label-smoothing side, where the original low-ranked boundary errors ✗ are less (rankwise) confident, such as the roof of the car and the right border of the median strip. Interestingly, LS-based high confidence errors in Figure 23 mainly correspond to a change of texture of the median strip – likely associated with an uncertainty on the true label – and correlate with those of the CE-based model. For brevity, we provide an image for a logit normalisation with norm dimension $p=5$, but the results do not differ much from the other values. As stated above, we will provide additional figures to readers upon request.
@@
Figure 22: Logit normalisation improves on boundary-related very confident errors ✗, but not on the bottom right errors ✗.
@@
Figure 23: Logit normalisation does not always fix selective classification when used on a model trained with a label-smoothing (right). However, boundary errors ✗ have lower (ranked) confidence, such as for the car’s roof.

### v3

@@ C.1 Image classification
For the DenseNet trained on CIFAR-100, we randomly split the original test set into 2000 images for validation and 8000 for evaluation. We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov (Nesterov, 1983; Sutskever et al., 2013), a momentum of 0.9, and $1\times 10^{-4}$ weight decay. We divide the learning rate by ten after 150 and 225 epochs. We use standard augmentations: we apply random crop with a four-pixel padding as well as random horizontal flip. We do not perform model selection and keep the last checkpoint.
@@
For our ViT-S-16 trained on ImageNet, we take batches of 2048 images and train on 4 A100s for 300 epochs with AdamW (Loshchilov & Hutter, 2019) with the $\beta$s equal to 0.9 and 0.999. We start with a linear warmup for 15 epochs, then use a cosine annealing scheduler with $3\times 10^{-3}$ as the starting learning rate. The models are trained with non-adaptive sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022) with $\rho=0.2$. We use a dropout (Srivastava et al., 2014) rate of 0.1 but no attention dropout. We transform the training images with a standard random resized crop to 224$\times$224 pixels using bicubic interpolation and a random horizontal flip. For evaluation, we center-crop the images to this resolution. For ImageNet, we do not perform model selection and keep the last checkpoint. However, we randomly extract a validation set of 10,000 images from the validation set to perform logit normalisation, and evaluation on the remaining 40,000.
@@
Our ResNet-50 is trained on 4 A100 with stochastic gradient descent for 120 epochs using a batch size of 1024 images. After five epochs of linear warmup, we use a cosine annealing scheduler starting with a learning rate of 0.4 with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We use the same transformation of the images as for the ViT and select the last checkpoint for inference. We use the same validation set as for the ViT for logit normalisation.
C.2 Semantic segmentation
@@
We train a Deeplabv3+ on CityScapes (Cordts et al., 2016) with a ResNet-101 (He et al., 2016) backbone pre-trained on ImageNet (Russakovsky et al., 2015). We use stochastic gradient descent with a base learning rate of 0.01, divided by 10 for the backbone weights, and reduced following the "poly" policy (Liu et al., 2015) with a power of 0.9. The weights are optimised with a momentum of 0.9 and a weight decay of $10^{-4}$. We take a batch size of 16 images and train for 30,000 steps. During training, we randomly crop the input images and targets to squares with 768-pixel-long sides. We apply random horizontal flip and colour-jitter with the classical parameters: brightness, contrast, and saturation levels of 0.5. For testing, we use the images at their original resolution and do not perform any test time augmentations. For the RC curves, we randomly sample 5000 predictions per image extracted prior to the final interpolation to compute the coverage and error rates. We keep the pixel-wise locations of the samples when changing the level of label smoothing $\alpha$ to ensure fair comparisons.
C.3 Tabular data
We perform the tabular data binary classification on two UCI datasets: Bank Marketing and Online shoppers. These two datasets have an input dimension and a number of samples of 7 and 45,211 for the former and 16 and 12,330 for the latter. In both cases, we select 80% of the data points for the test set and train the models for 10 epochs using PyTorch’s default Adam optimizer and a batch size of 128. Similarly to the other experiments of this paper, we do not perform model selection and keep the last checkpoint. Our models are MLPs with two hidden layers, whose size depends on the input size as follows: the first hidden layer has 4 times as many neurons as the dimension of the input data, and the second has half the number of neurons of the first layer.
C.4 Natural language data
For the training on the IMDB (Maas et al., 2011) dataset, we tokenize the data with nltk and convert them to 300-dimensional vectors using GloVe 840B (Pennington et al., 2014). The architecture of the models is defined as follows:
  • an LSTM layer with a 300 input dimension and a 256 output dimension, followed by a dropout layer of rate 0.2,
  • a linear layer keeping the dimension of the vectors unchanged, on which we apply ReLU, followed by a second dropout layer with the same rate,
  • two linear layers with ReLU reducing the dimension to 128 and 1, respectively (binary classification).
Finally, we train the models with PyTorch’s Adam default for 10 epochs using the pre-made train test split. The code is available on GitHub.
C.5 Logit Calculations
We perform logit evaluation calculations in double precision (softmax, logit normalisation, entropy etc.), in order to reduce the impact of numerical error.
@@ C.5 Logit Calculations
Appendix D Analysis of empirical loss (one-hot labels)
@@
which in turn gives the suppression gradient,
$\frac{\partial\mathcal{L}_{\text{sup}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}-\mathcal{L}_{\text{CE}})}{\partial v_{k}}=\frac{\partial\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\partial\mathcal{L}_{\text{CE}}}{\partial v_{k}}=\alpha\left[\delta_{y\omega_{k}}-1/K\right]~{},$ (20)
which is once again independent of the model output $\boldsymbol{\pi}$. In this case, as the label $y$ has already been sampled, the model is either right  or wrong , giving two different suppression gradients for the max logit $v_{\text{max}}$,
@@
We can see that the max logit is more strongly suppressed during training when the prediction is correct, which aligns with the analysis of $\mathcal{L}^{\text{true}}$ and $P_{\text{error}}$ in the main paper. When the model is correct during training, the max logit is suppressed, but when it is incorrect the max logit is not suppressed. Thus the uncertainty ranking of correct ✓ vs incorrect ✗ data samples is degraded, harming SC. This leads us to the same conclusions as in Sec. 4 and also aligns with the behaviour in Fig. 5 (right) where $v_{\text{max}}$ is higher for errors given the value of MSP.
Appendix E Result and proof
@@
Similar to Eq. 23, please note that using $k=i$, we get the same equation as Eq. 25, although with a strict inequality. We can now sum on the elements of $\boldsymbol{v}$ to get
@@
By taking the inverse (all values are strictly positive), we get
@@
Appendix F Discussions
F.1 $U$ other than MSP
@@
Finally, we also choose to omit various training/architecture-based approaches such as (Ziyin et al., 2019; Geifman & El-Yaniv, 2019; Moon et al., 2020; Zhu et al., 2024), as we aim to focus our investigation solely on understanding the training effects of LS. We note that LS (potentially combined with post-hoc logit normalisation) is considerably simpler to implement than the aforementioned training-based approaches, and has been shown to be stable in many more use cases and so may be more likely to be chosen by a practitioner. Besides, recent work (Feng et al., 2023) has suggested that MSP applied to the methods in (Ziyin et al., 2019; Geifman & El-Yaniv, 2019) actually performs better than the $U$s proposed in them (reject logit/selection head).
@@
F.2 Aleatoric and epistemic uncertainty
@@
F.3 On the assumption that logits are positive
@@
Largest positive logits dominate.
Result 1 assumes that all elements in the logit vector are $>0$, which is not necessarily true. However, we also find empirically that both $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ tend to be dominated by the largest positive logits. This is intuitive as exponentiating or raising to power $p>1$ will amplify the larger logits. This is shown in Fig. 21, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set ($p=5$ is optimal on the validation data for logit normalisation in this case). Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive and dominate the computation. As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to increase uncertainty for higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$, even if not all $v_{k}>0$. (Note that we omit shift $s=-1/K\sum_{k}v_{k}$ as it is $\approx 0$ in our experiments).
@@
Logit centralisation.
We follow Cattelan & Silva (2024) and suggest to shift the logits by their mean during normalisation,
$\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}-\mu(\boldsymbol{v})}{\left\lVert\boldsymbol{v}-\mu(\boldsymbol{v})\right\rVert_{p}},\quad\mu(\boldsymbol{v})=\frac{1}{K}\sum_{k}v_{k}.$ (28)
In our experiments, this has little-to-no effect as $\mu\approx 0$ (for example for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation data $\mu$ averages $\sim 3\times 10^{-3}$ with std. $\sim 1\times 10^{-3}$). This is corroborated by Cattelan & Silva (2024)’s observations in their Appendix G, where they find this shift does not affect the vast majority of their models.
We note that according to the analysis in Sec. 5.1 and Sec. F.3, for logit normalisation to be effective, $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ should be dominated by positive logits. Logits are not necessarily constrained to be positive. They can be arbitrarily shifted without affecting the cross entropy loss (e.g. via bias $\boldsymbol{b}$ of the final layer), or derived as log probabilities which are necessarily negative. Thus we suggest applying the above shift to ensure that the largest logits are positive. We note that other shifts are possible (for example $s=-\min_{k}v_{k}$) to achieve the same purpose but we leave the exploration of this to future work.
F.4 Explanations of Logit normalisation in Cattelan & Silva (2024)
The work that introduces logit normalisation (Cattelan & Silva, 2024) is primarily an experimental study, where the focus is on extracting empirical takeaways. However, the authors do discuss potential reasons for the effectiveness of the approach in their Appendix B. They observe that models that are generally (on both ✓ and ✗) more uncertain benefit from logit normalisation and so suggest that logit normalisation alleviates “underconfidence” by reducing the influence of lower-ranked logits on the uncertainty of a prediction (using analysis similar to ours in Appendix E). We believe their explanation is ultimately incomplete as:
@@ F.4 Explanations of Logit normalisation in Cattelan & Silva (2024)
    They do not clearly delineate the definition of over/underconfidence used in model calibration with that used in selective classification. This is an issue since calibration is concerned with absolute marginal (averaged over data samples) properties, whilst selective classification is concerned with relative conditional (per sample) properties. It is possible to be extremely over/underconfident in the calibration sense, whilst being optimal for SC, or very well calibrated and worse at distinguishing correct vs incorrect samples (Zhu et al., 2024).
@@
F.5 Analysis of Negative Label Smoothing (Wei et al., 2022)
We consider the framework developed in the main paper to study Generalized Label Smoothing (GLS) with negative smoothing values – called Negative Label Smoothing (NLS) – as suggested by Wei et al. (2022). The definition of GLS is the same as the original from Szegedy et al. (2016), except for the domain of the label smoothing value $\alpha$, which can take values in $(-\infty,1]$. Wei et al. (2022) suggest that using negative values for the label smoothing value can lead to improved performance when training with noisy labels as well as in the “clean” setting. It is tempting to consider, as negative $\alpha$ ought to reverse the logit suppression described in Eq. 15. However, here we show that NLS can lead to unstable training characteristics.
The per-sample cross-entropy between a single softmax prediction $\pi$ and the corresponding true categorical distribution $\bar{\pi}$ is convex and can be written as follows:
$\text{CE}(\pi,\bar{\pi})=-\sum_{k\in\llbracket 1,K\rrbracket}\left[(1-\alpha)\bar{\pi}_{k}+\frac{\alpha}{K}\boldsymbol{1}\right]\log\pi_{k}.$ (29)
For $\alpha\geq 0$, which covers vanilla cross entropy ($\alpha=0$) and regular label smoothing ($\alpha>0$) a single global minimum is attained in $\pi=(1-\alpha)\bar{\pi}+\frac{\alpha}{K}\boldsymbol{1}$. Here, the gradient of the loss is zero therefore creating a balance and stabilizing training, taking the gradient as in Eq. 13,
${\color[rgb]{0,0,0}\frac{\partial\text{CE}(\pi,\bar{\pi})}{\partial v_{k}}=-\biggl{[}\underbrace{\Bigl{[}(1-\alpha)\bar{\pi}_{k}+\alpha/K\Bigr{]}}_{\text{target}}\!-\underbrace{\pi_{k}}_{\text{softmax output}}\biggr{]}.}$ (30)
However, in the case of a negative label smoothing value $\alpha<0$ the target can be outside $[0,1]$, but the softmax output is constrained to $[0,1]$, thus $\pi=(1-\alpha)\bar{\pi}+\frac{\alpha}{K}\boldsymbol{1}$ is not achievable and the gradient will always have some magnitude and there is no optimisation minimum.
Indeed, if we consider for one-hot $\bar{\pi}$, the $i$-th terms s.t. $\bar{\pi}_{i}=0$ of Eq. 29 will be strictly negative with a negative-label-smoothed target. Since the log of the softmax probabilities $\pi$ can take values in $(-\infty,0]$, so can the loss. Being convex, there cannot exist a local minimum, and therefore, the gradient can never be zero and pushes its value to $\infty$- ∞ by reducing the value of the corresponding logit to $\infty$- ∞ and pushing the others to $\infty$. We hypothesise that this unbalance explains the instability of the method that users reported on the GitHub repository of the original paper. Unfortunately, we too were not able to reproduce the experiments of Wei et al. (2022) and obtain stable enough training runs to perform experiments on NLS and confirm this theoretical analysis.
F.6 Existing benchmarks and training recipes with LS
Appendix G Additional segmentation results
In this section, we complete the picture by providing additional semantic segmentation results to Figs. 1 and 8. As in the main paper, the following segmentation maps are performed by ResNet-101-based DeepLab-v3+ trained on Cityscapes with cross-entropy and label-smoothing with $\alpha=0.2$. We recall that the models and notebooks used to generate these plots are available on Hugging Face and GitHub, respectively. In these figures, we provide the rank of all the predictions – not only errors – to show the difference between the model’s behavior on its errors but also correct predictions. The rank of the correct predictions ✓ is slightly greyed out compared to the rank of the errors ✗. Fig. 23 presents the full ranks corresponding to the scene used in the main paper.
@@ Appendix G Additional segmentation results
Figs. 23 and 24 shows that the cross-entropy-based model predicts label maps in which pixel uncertainty ranks are generally smoother than those provided by the label-smoothing-based model. The label-smoothing model exhibits large areas with low uncertainty rankings, which suddenly increases near the boundary of objects. However, the boundaries are sometimes wrongly predicted, leading to (rank-wise) highly confident errors ✗. The pixels corresponding to the boundaries of objects, where the probability of error is naturally higher and which are most sensitive to (ground-truth) label errors, have higher confidence (lower uncertainty ranks) according to the LS-based model than the CE-based model. This explains the high-confidence errors ✗ in the misclassified boundaries of several objects, either in Figs. 1 and 24, and aligns with the imbalanced suppression of Eq. 15. We again see how logit normalisation is able to decrease the confidence of the label-smoothing-trained model on highly confident (bright yellow) errors ✗.
[…]
@@
Figure 24: Logit normalisation improves on boundary-related very confident errors ✗, but does not completely fix the high (ranked) confidence of the bottom right errors ✗.
@@
Figure 25: Logit normalisation does not always completely fix selective classification when used on a model trained with a label-smoothing (right).
Appendix H Impact and Future Work
In this section, we provide some additional discussion about the (practical) impact of our work as well as promising directions for potential future research.
Direct impact.
By empirically verifying and analytically elucidating the limited experimental results pertaining to LS in (Zhu et al., 2024) (it is not the focus of that work), we provide strong evidence that the behaviour that LS degrades SC is generalisable (over architectures, data modalities etc.). In particular, we directly analyse the loss, which is common to all settings involving label smoothing. This will help inform practitioners of selective classification when they are designing and deploying systems. Furthermore, by explaining the efficacy of logit normalisation we provide an effective and well-motivated solution to the previously demonstrated problem. We emphasise that when logit normalisation was introduced in (Cattelan & Silva, 2024), it was not clearly explained why logit normalisation was effective on some pretrained models and ineffective on others. In our work, we analytically clarify the mechanism of logit normalisation, opening up the figurative black box, and are able to directly link it to our previous analysis on LS. This provides clear guidance on when and why to use logit-normalisation, giving potential practitioners confidence in the effectiveness of the approach, which is especially important in high-risk safety-critical applications.
Future work.
The empirical and analytical results relating to LS in our work naturally suggest that other training approaches that alter the labels such as Mixup (Zhang et al., 2018) may also have similar adverse effects and/or be amenable to similar gradient analysis. It also raises the question of how such label augmentations effect problem settings outside of selective classification such as OOD detection or transfer learning. The novel analysis of LS in Sec. 4 may also be useful in problem settings beyond SC: Does this aspect of label smoothing (suppressing the max logit less for incorrect predictions) affect generalisation? What about behaviour on OOD data? Could it help explain the behaviour in (Kornblith et al., 2021) where LS results in worse transfer/representation learning? Can this insight lead to a modification of LS to improve it? Our analysis of the mechanism of logit normalisation in Sec. 5.1 is also general – we simply demonstrate that logit normalisation reduces confidence when the max logit is higher. Thus, this knowledge can be potentially applied to other scenarios (e.g. OOD detection). Furthermore, in the case of post-hoc methods, logit normalisation using the $p$-norm is just one possible option for correcting the effect of LS – armed with the insight presented in our work it may be possible to develop superior or more principled methods for extracting better uncertainty estimates from logits.